OpenStudy (anonymous):

what is 3x + y = 2 in Y=mx+b form? im too lazy -_-

OpenStudy (anonymous):

rearrange for y=mx+b 3x+y = 2 bring the y to the right hand side and 2 to the left hand side. 3x-2=-y -3x+2=y y=-3x+2

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

when graphing it would be a negative slope right?

OpenStudy (anonymous):

yes u are right because the gradient(m) is -3.

OpenStudy (anonymous):

if the gradient positive it goes upwards, downwards if it is negative.

OpenStudy (anonymous):

thats going down 3, left 1 or down 3, right one

OpenStudy (anonymous):

um no

OpenStudy (anonymous):

my bad.

OpenStudy (anonymous):

\[\frac{rise}{run}=\frac{-3}{1}\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

rise run we get from any two points on the line.

OpenStudy (anonymous):

I'm sure u know that?

OpenStudy (anonymous):

heh yeah...... duh? who does'nt?

OpenStudy (anonymous):

haha kk.

OpenStudy (anonymous):

the gradient can also be calculated by using the gradient formula. (y2-y1) / (x2-x1) say we have two points on the straight line (5,1) (2,4) x1,y1 x2,y2 using the gradient formula m=(4-1) / (2-5) m=3/-3 m=-1

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

we did that in class, i just did'nt know it was called that